OpenStudy (anonymous):

Find the derivative of k(x)=4/x by using the power rule. k ' (x) = __________ how do we start this??

OpenStudy (agent0smith):

*exponent laws

OpenStudy (agent0smith):

\[\Large k(x) = 4 x^{-1}\]

OpenStudy (anonymous):

oh okay:) so to make it a derivative what happens?

OpenStudy (agent0smith):

Use the power rule and find out :P multiply by the power, then reduce the power by one

OpenStudy (anonymous):

so 4x^-1 * -1 ?

OpenStudy (agent0smith):

You missed something, read the power rule again

OpenStudy (anonymous):

so k(x)=4/x = 4x^-1 f ' (x) = nx^n-1 so 4x^4-1 = 4x^3 ?

OpenStudy (agent0smith):

You're getting things mixed up. The power is -1, not 4.

OpenStudy (anonymous):

ahh okay :/ erm so k(x) simplifies to 4x^-1 right?

OpenStudy (anonymous):

and the rule is f ' (x) = nx^(n-1) right? :/

OpenStudy (agent0smith):

\[\Large k(x) = 4 x^{-1} \]multiply by the power (which is the -1). Then reduce the power by 1.

OpenStudy (anonymous):

=4(-x) ?

OpenStudy (agent0smith):

\[\Large k'(x) = (-1)*4 x^{-1-1}\]

OpenStudy (anonymous):

ohhh okay... soo, (-1) * 4x^-2 ? so (-1) * 4x = -4x ?

OpenStudy (agent0smith):

I can not figure out what you are doing...

OpenStudy (anonymous):

ahh okay sorry.. umm i messed up haha so (-1) * 4x^-2 = -4x^-2 = -4 / x^2 ??

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

ah okay awesome!! thank you!!

OpenStudy (agent0smith):

:)